# Orbifold modifications of complex analytic varieties

### v4

@@ 7. Modification around an irreducible singular locus
Then the finite morphism $\overline{V}_{k}\to\iota_{k}(V)$ is constructed by applying Proposition 6.7 on $\zeta_{k}(Z)\subseteq X_{\sigma(k)}$, and $\overline{V}_{l}\to\iota_{l}(V_{l})$ is constructed by applying Proposition 6.7 on $\zeta_{l}(Z)\subseteq X_{\sigma(l)}$. By the property (5) of Proposition 6.7, we deduce that, up to shrinking $U$, every connected component of $\overline{V}_{k}$ is isomorphic to every connected component of $\overline{V}_{l}$. Lemma 2.11 then implies that $(\overline{W}_{k},\overline{G}_{k})_{k\in K}$ induces a complex analytic orbispace $\overline{\mathfrak{W}}$. Its quotient space is exactly $W^{\prime}$ since $W^{\prime}$ and every $\overline{W}_{k}$ are normal. Replacing $Y$ by its normalization, we can assume that $W^{\prime}=W$, which is an open neighborhood of $S_{Y}$ in $Y$.
It remains to prove the items (4) and (5). We notice that the divisorial critical locus of $W_{k}\to W$ is contained in $\Delta_{Y}$, as the divisorial critical locus of $X_{\sigma(k)}\to X$ is contained in $\Delta$. Since the divisorial locus of $\overline{W}_{k}\to W_{k}$ is contained in the preimage of $\Delta_{Y}$ by construction, we deduce the item (4). The item (5) follows from the item (3) of Proposition 6.7. This completes the proof of the lemma. ∎
@@
Let $\mathfrak{X}=(X_{k},G_{k})_{k\in K}$ be the complex analytic orbispace induced by the index-one covers of $\omega_{X}$. Then $X_{k}\to X$ is quasi-étale onto its image, and $X_{k}$ has at most canonical singularities. Let $S_{k}$ be the preimage of $S$ in $X_{k}$. We note that $S_{k}$ may not be irreducible. Let $h_{k}\colon X^{\prime}_{k}\to X_{k}$ be the bimeromorphic projective morphism induced by the functorial desingularization of $S_{k}$, see the beginning of Subsection 3.B. By construction, the strict trasnform $S^{\prime}_{k}=(h_{k}^{-1})_{*}S_{k}$ is smooth. There is a natural action of $G_{k}$ on $X^{\prime}_{k}$ and we obtain a complex analytic orbispace $\mathfrak{X}^{\prime}=(X^{\prime}_{k},G_{k},\mu_{k})_{k\in K}$ with quotient space $X^{\prime}$. Furthermore, there is a natural projective bimeromorphic morphism $h\colon X^{\prime}\to X$, with a $h$-ample and $h$-exceptional divisor. We remark that the indeterminacy locus of $h^{-1}$ is a proper closed subset of $S$. Moreover, the divisorial critical locus of $\mu_{k}$ is contained in the $h$-exceptional locus.
First we assume that $X_{k}$ is smooth at general points of the components of $S_{k}$. Then $X^{\prime}_{k}$ is smooth around general points of the components of $S^{\prime}_{k}$. Let $Y_{k}\to X^{\prime}_{k}$ be the projective bimeromorphic morphism as in Corollary 4.6, with respect to $S^{\prime}_{k}$. If $(S_{Y})_{k}$ is the strict transform of $S^{\prime}_{k}$ in $Y_{k}$, then $Y_{k}$ is smooth around $(S_{Y})_{k}$. Up to taking the normalization, we can assume that every $Y_{k}$ is normal. Since the constructions are functorial, there is a natural action of $G_{k}$ on $Y_{k}$, and we have a complex analytic orbispace $\mathfrak{Y}=(Y_{k},G_{k},\psi_{k})_{k\in K}$ with quotient space $Y$, which is normal. Furthermore, there is a natural projective bimeromorphic morphism $f\colon Y\to X$. Since $Y_{k}$ is smooth around $(S_{Y})_{k}$, there is a smooth $G_{k}$-invariant open neighborhood $W_{k}$ of $(S_{Y})_{k}$. Then $(W_{k},G_{k},\psi_{k})_{k\in K}$ is a complex orbifold, whose quotient space is an open neighborhood $W$ of $S_{Y}$. Hence $f$ satisfies the properties of the proposition.
@@
We apply Lemma 7.3 on $Z$ and obtain a projective bimeromorphic morphism $h_{1}\colon Z_{1}\to Z$. Since the indeterminacy locus of $h_{1}^{-1}$ is contained in $S^{\prime}$, which is compact, we deduce from Lemma 2.3 that $h_{1}$ extends to a proper bimeromorphic morphism $Y_{1}\to X^{\prime}$. Let $f_{1}\colon Y_{1}\to X$ be the natural morphism. Then there is a $f_{1}$-ample and $f_{1}$-exceptional divisor. Let $S_{Y_{1}}$ be the strict transform of $S^{\prime}$ in $Y_{1}$, and let $H_{Y_{1}}$ be the reduced divisor in $S_{Y_{1}}$ as in the item (3) of Lemma 7.3. By applying Lemma 7.4 around $S_{Y_{1}}$, we can obtain an open neighborhood $Q_{1}$ of it, so that we can apply Lemma 7.5 to $Q_{1}$. Then we obtain a projective bimeromorphic morphism $g_{2}\colon P_{2}\to Q_{1}$. By Lemma 2.3 again, $g_{2}$ extends to a projective bimeromorphic morphism $g_{2}\colon Y_{2}\to Y_{1}$. Let $f_{2}\colon Y_{2}\to X$ be the induced morphism.
@@
We can apply Lemma 7.6 to $Q_{2}$ with the divisor $\Gamma_{2}\cap Q_{2}=\Delta_{2}\cap Q_{2}$, and obtain a projective bimeromorphic morphism $g_{3}\colon P_{3}\to Q_{2}$. By Lemma 2.3, we deduce that $g_{3}$ extends to a projective bimeromorphic morphism $g_{3}\colon Y_{3}\to Y_{2}$. Let $Y=Y_{3}$ and let $f\colon Y\to X$ be the natural morphism. Then $f$ satisfies all the properties of the proposition. ∎
@@ 8. Proof of the modification theorem
    The divisorial critical locus of $\psi_{i}\colon W_{i}\to W$ is contained in $\Delta$.
@@
Assume that there is a projective bimeromorphic morphism $h\colon(Z,\Gamma)\to(X,\Delta)$ satisfying the following properties. The divisor $\Gamma$ is the sum of $h^{-1}_{*}\Delta$ and the whole reduced $h$-exceptional divisor. The $h$-exceptional locus is contained in $\Gamma$. The pair $(Z,\Gamma)$ is dlt and each irreducible component of $\Gamma$ is $\mathbb{Q}$-Cartier. The morphism $h$ is an isomorphism over general points of $S$. The pair $(Z,\Gamma)$ is snc on $h^{-1}(X\setminus S)$.
Then there is a projective bimeromorphic morphism $\varphi\colon Y\to Z$ with a $\varphi$-ample and $\varphi$-exceptional divisor, such that $Y$ has quotient singularities. Furthermore, $\varphi$ is an isomorphism over $h^{-1}(X\setminus S)$ and over general points of the strict transform of $S$ in $Z$.
@@
[…]
Since $V$ is normal, the family $(V_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace whose quotient space is $V$. Since $Z\setminus\Gamma$ is isomorphic to an open subset of $X\setminus\Delta$, we deduce from the condition (2) that, if $\mu_{i}\colon V_{i}\to V$ is the natural morphism, then its divisorial critical locus is contained in $\Gamma$. In addition, the condition (3) implies that $V_{i}$ is smooth on $\mu_{i}^{-1}(V\setminus\Gamma)$.
Let $S_{Z}$ be the strict transform of $S$ in $Z$. Let $\Gamma^{j}$ with $j=1,...,l$ be the irreducible components of $\Gamma$, and let $r^{j}$ be the Cartier index of $\Gamma^{j}$. For every point $z\in h^{-1}(S)$, there is an open neighborhood $T$ of $z$ in $V$ such that each $r^{j}\Gamma^{j}|_{T}$ is defined by some holomorphic function $\sigma^{j}$. Thus we obtain a family $(U_{k})_{k\in K}$ of open subsets of $V$, whose union contains $h^{-1}(S)$, such that $r^{j}\Gamma^{j}|_{U_{k}}$ is connected and defined by some holomorphic function $\sigma^{j}_{k}$. Furthermore, there is an application $\iota\colon K\to I$ such that $U_{k}$ is contained in $\mu_{\iota(k)}(V_{\iota(k)})$. By convention, we set $\sigma_{k}^{j}=1$ if $\Gamma^{j}|_{U_{k}}$ is empty. Let $U=\bigcup_{k\in K}U_{k}$. Then it is a neighborhood of $h^{-1}(S)$ in $V$.
@@
[…]
and we let $\xi_{k}\colon\overline{U}_{k}\to U_{k}$ be the natural morphism. Then the $\overline{U}_{k}$’s define a normal complex analytic orbispace with quotient space $U$, by Lemma 2.6 and Lemma 2.11. Let $Q_{k}$ be a connected component of the normalization of $\overline{U}_{k}\times_{Z}V_{\iota(k)}$. It follows that the natural finite morphism $Q_{k}\to U_{k}$ is Galois of Galois group $H_{k}$, and the collection $(Q_{k},H_{k})_{k\in K}$ induces a normal complex analytic orbispace $\mathfrak{Q}$ with quotient space $U$.
[…]
We denote by $\eta_{k}\colon Q_{k}\to\overline{U}_{k}$ the natural morphism. Then by Lemma 2.8, we obtain that $\eta_{k}$ is quasi-étale; furthermore, $Q_{k}$ is smooth in codimension 2 since $V_{\iota(k)}$ is smooth over $U_{k}\setminus\Gamma$.
@@
Let $\widetilde{Q}_{k}\to Q_{k}$ be the functorial desingularization, which is an isomorphism over the smooth locus of $Q_{k}$. Then there is a natural action of $H_{k}$ on $\widetilde{Q}_{k}$, and we obtain a complex orbifold $\widetilde{\mathfrak{Q}}=(\widetilde{Q}_{k},H_{k})_{k\in K}$ with quotient space $\widetilde{Q}$. In particular, $\widetilde{Q}$ has quotient singularities only. There is an induced projective bimeromorphic morphism $\varphi\colon\widetilde{Q}\to U$. From the properties of $(Q_{k})_{\operatorname{\textsubscript{\rm sing}}}$ above, we see that the indeterminacy locus of $\varphi^{-1}$ is contained in $h^{-1}(S)$ and is of codimension at least 3 in $Z$. In particular, $\varphi$ is an isomorphism over general points of $S_{Z}$. Furthermore, there is an $\varphi$-exceptional and $\varphi$-ample divisor. Thus by Lemma 2.3, $\varphi$ extends to a projective bimeromorphic morphism $\varphi\colon Y\to Z$. Since $Z$ is smooth outside $h^{-1}(S)$, we deduce that $Y$ has quotient singularities only. This completes the proof of the lemma. ∎
@@
We will divide the proof into several steps. Let $S_{e}$ with $e=1,...,r$ be all the irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$ which have codimension 2 in $X$.
@@
Step 1. In this step, we will reduce to the case when $S_{1},...,S_{r}$ are pairwise disjoint. By blowing up at centers contained in at least two of the $S_{e}$’s, we obtain a projective bimeromorphic morphism $h\colon Z\to X$ such that the strict transforms $h^{-1}_{*}S_{1},...,h^{-1}_{*}S_{r}$ are pairwise disjoint. Let $X^{\prime}\to Z$ be the projective bimeromorphic morphism constructed in Lemma 2.18, with respect to the collection $\{h^{-1}_{*}S_{1},...,h^{-1}_{*}S_{r}\}$. We observe that the indeterminacy locus of $X\dashrightarrow X^{\prime}$ has codimension at least 3. Up to replacing $X$ by $X^{\prime}$, we may assume that $S_{1},...,S_{r}$ are pairwise disjoint. Furthermore, $X$ has klt singularities.
@@
Step 2. In this step, we will reduce to the situation of Lemma 8.1. Let $X_{e}$ be an open neighborhood of $S_{e}$ in $X$ such that the $X_{e}\cap X_{e^{\prime}}=\emptyset$ if $e\neq e^{\prime}$. We apply Proposition 7.1 to each $X_{e}$. Then we have projective bimeromorphic morphisms $p_{e}\colon X^{\prime}_{e}\to X_{e}$, which satisfies the properties of Proposition 7.1. Since for each $e$, there is a $p_{e}$-ample and $p_{e}$-exceptional divisor, and since the indeterminacy locus of $(p_{e})^{-1}$ is contained in $S_{e}$ and hence compact, by applying Lemma 2.3 for several times, there is a projective bimeromorphic morphism $p\colon X^{\prime}\to X$ such that $p=p_{e}$ over $X_{e}$. Let $\Delta$ be the whole reduced $p$-exceptional divisor. Then $p(\Delta)$ has codimension at least 3 in $X$.
@@
In the remainder of the proof, by abuse of notation, we denote $X^{\prime}$ by $X$, and $X_{e}^{\prime}$ by $X_{e}$. If $\Delta_{e}=\Delta|_{X_{e}}$, then $X_{e}$, $\Delta_{e}$ and $S_{e}$ satisfy the conditions (1)-(3) of Lemma 8.1. Furthermore, from now on, it is enough to construct $f\colon Y\to X$, such that $Y$ has quotient singularities, and $f$ is an isomorphism over general points of each $S_{e}$ and over $X_{\operatorname{\textsubscript{\rm sm}}}\setminus\Delta$. We remark that $X_{\operatorname{\textsubscript{\rm sm}}}\setminus\Delta$ is contained in the snc locus of $(X,\Delta)$.
@@
Step 3. In this step, we will construct a bimeromorphic morphism $h\colon Z\to X$, which plays the same role for $X_{e}$ as $Z$ for $X$ in Lemma 8.1. We first note that, by the property (3) of Lemma 8.1, $X$ has quotient singularities around $S\setminus\Delta$, where $S=\bigcup_{e=1}^{r}S_{e}$. Thus, we can apply Corollary 3.4 to $(X,\Delta)$, with respect to $S$, and obtain a projective bimeromorphic morphism $\rho\colon X^{\prime}\to X$ with $X^{\prime}$ normal. By construction, $\rho$ is an isomorphism over the snc locus of $(X,\Delta)$ and over general points of each $S_{e}$; if $\Delta^{\prime}$ is the sum of $\rho^{-1}_{*}\Delta$ and the whole reduced $\rho$-exceptional divisor, then $(X^{\prime},\Delta^{\prime})$ is snc on $\rho^{-1}(X\setminus S)$; furthermore, $\Delta^{\prime}$ is the support of some Cartier divisor, and $X^{\prime}\setminus\Delta^{\prime}$ has quotient singularities. Let $S^{\prime}$ be the strict transform of $S$ in $X^{\prime}$. Let $(Z,\Gamma)\to(X^{\prime},\Delta^{\prime})$ be a $\mathbb{Q}$-factorial dlt modification as in Lemma 2.19, which is an isomorphism over the snc locus of $(X^{\prime},\Delta^{\prime})$, and over general points of each component of $S^{\prime}$. We denote by $h\colon Z\to X$ the natural morphism. Then we have the following properties. The divisor $\Gamma$ is the sum of $h^{-1}_{*}\Delta$ and the whole reduced $h$-exceptional divisor. The $h$-exceptional locus is contained in $\Gamma$. Each irreducible component of $\Gamma$ is $\mathbb{Q}$-Cartier. The morphism $h$ is an isomorphism over the snc locus of $(X,\Delta)$ and over general points of each component of $S$. The pair $(Z,\Gamma)$ is snc on $h^{-1}(X\setminus S)$.

### v6

@@ 7. Modification around an irreducible singular locus
Then the finite morphism $\overline{V}_{k}\rightarrow\iota_{k}(V)$ is constructed by applying Proposition 6.7 on $\zeta_{k}(Z)\subseteq X_{\sigma(k)}$, and $\overline{V}_{l}\rightarrow\iota_{l}(V_{l})$ is constructed by applying Proposition 6.7 on $\zeta_{l}(Z)\subseteq X_{\sigma(l)}$. By the property (5) of Proposition 6.7, we deduce that, up to shrinking $U$, every connected component of $\overline{V}_{k}$ is isomorphic to every connected component of $\overline{V}_{l}$. Lemma 2.11 then implies that $(\overline{W}_{k},\overline{G}_{k})_{k\in K}$ induces a complex analytic orbispace $\overline{\mathfrak{W}}$. Its quotient space is exactly $W^{\prime}$ since $W^{\prime}$ and every $\overline{W}_{k}$ are normal. Replacing $Y$ by its normalization, we can assume that $W^{\prime}=W$, which is an open neighborhood of $S_{Y}$ in $Y$.
It remains to prove the items (4) and (5). We notice that the divisorial critical locus of $W_{k}\rightarrow W$ is contained in $\Delta_{Y}$, as the divisorial critical locus of $X_{\sigma(k)}\rightarrow X$ is contained in $\Delta$. Since the divisorial locus of $\overline{W}_{k}\rightarrow W_{k}$ is contained in the preimage of $\Delta_{Y}$ by construction, we deduce the item (4). The item (5) follows from the item (3) of Proposition 6.7. This completes the proof of the lemma. ∎
@@
Let $\mathfrak{X}=(X_{k},G_{k})_{k\in K}$ be the complex analytic orbispace induced by the index-one covers of $\omega_{X}$. Then $X_{k}\rightarrow X$ is quasi-étale onto its image, and $X_{k}$ has at most canonical singularities. Let $S_{k}$ be the preimage of $S$ in $X_{k}$. We note that $S_{k}$ may not be irreducible. Let $h_{k}\colon X^{\prime}_{k}\rightarrow X_{k}$ be the bimeromorphic projective morphism induced by the functorial desingularization of $S_{k}$, see the beginning of Subsection 3.B. By construction, the strict trasnform $S^{\prime}_{k}=(h_{k}^{-1})_{*}S_{k}$ is smooth. There is a natural action of $G_{k}$ on $X^{\prime}_{k}$ and we obtain a complex analytic orbispace $\mathfrak{X}^{\prime}=(X^{\prime}_{k},G_{k},\mu_{k})_{k\in K}$ with quotient space $X^{\prime}$. Furthermore, there is a natural projective bimeromorphic morphism $h\colon X^{\prime}\rightarrow X$, with a $h$-ample and $h$-exceptional divisor. We remark that the indeterminacy locus of $h^{-1}$ is a proper closed subset of $S$. Moreover, the divisorial critical locus of $\mu_{k}$ is contained in the $h$-exceptional locus.
First we assume that $X_{k}$ is smooth at general points of the components of $S_{k}$. Then $X^{\prime}_{k}$ is smooth around general points of the components of $S^{\prime}_{k}$. Let $Y_{k}\rightarrow X^{\prime}_{k}$ be the projective bimeromorphic morphism as in Corollary 4.6, with respect to $S^{\prime}_{k}$. If $(S_{Y})_{k}$ is the strict transform of $S^{\prime}_{k}$ in $Y_{k}$, then $Y_{k}$ is smooth around $(S_{Y})_{k}$. Up to taking the normalization, we can assume that every $Y_{k}$ is normal. Since the constructions are functorial, there is a natural action of $G_{k}$ on $Y_{k}$, and we have a complex analytic orbispace $\mathfrak{Y}=(Y_{k},G_{k},\psi_{k})_{k\in K}$ with quotient space $Y$, which is normal. Furthermore, there is a natural projective bimeromorphic morphism $f\colon Y\rightarrow X$. Since $Y_{k}$ is smooth around $(S_{Y})_{k}$, there is a smooth $G_{k}$-invariant open neighborhood $W_{k}$ of $(S_{Y})_{k}$. Then $(W_{k},G_{k},\psi_{k})_{k\in K}$ is a complex orbifold, whose quotient space is an open neighborhood $W$ of $S_{Y}$. Hence $f$ satisfies the properties of the proposition.
@@
We apply Lemma 7.3 on $Z$ and obtain a projective bimeromorphic morphism $h_{1}\colon Z_{1}\rightarrow Z$. Since the indeterminacy locus of $h_{1}^{-1}$ is contained in $S^{\prime}$, which is compact, we deduce from Lemma 2.3 that $h_{1}$ extends to a proper bimeromorphic morphism $Y_{1}\rightarrow X^{\prime}$. Let $f_{1}\colon Y_{1}\rightarrow X$ be the natural morphism. Then there is a $f_{1}$-ample and $f_{1}$-exceptional divisor. Let $S_{Y_{1}}$ be the strict transform of $S^{\prime}$ in $Y_{1}$, and let $H_{Y_{1}}$ be the reduced divisor in $S_{Y_{1}}$ as in the item (3) of Lemma 7.3. By applying Lemma 7.4 around $S_{Y_{1}}$, we can obtain an open neighborhood $Q_{1}$ of it, so that we can apply Lemma 7.5 to $Q_{1}$. Then we obtain a projective bimeromorphic morphism $g_{2}\colon P_{2}\rightarrow Q_{1}$. By Lemma 2.3 again, $g_{2}$ extends to a projective bimeromorphic morphism $g_{2}\colon Y_{2}\rightarrow Y_{1}$. Let $f_{2}\colon Y_{2}\rightarrow X$ be the induced morphism.
@@
We can apply Lemma 7.6 to $Q_{2}$ with the divisor $\Gamma_{2}\cap Q_{2}=\Delta_{2}\cap Q_{2}$, and obtain a projective bimeromorphic morphism $g_{3}\colon P_{3}\rightarrow Q_{2}$. By Lemma 2.3, we deduce that $g_{3}$ extends to a projective bimeromorphic morphism $g_{3}\colon Y_{3}\rightarrow Y_{2}$. Let $Y=Y_{3}$ and let $f\colon Y\rightarrow X$ be the natural morphism. Then $f$ satisfies all the properties of the proposition. ∎
@@ 8. Proof of the modification theorem
    The divisorial critical locus of $\psi_{i}\colon W_{i}\rightarrow W$ is contained in $\Delta$.
@@
Assume that there is a projective bimeromorphic morphism $h\colon(Z,\Gamma)\rightarrow(X,\Delta)$ satisfying the following properties. The divisor $\Gamma$ is the sum of $h^{-1}_{*}\Delta$ and the whole reduced $h$-exceptional divisor. The $h$-exceptional locus is contained in $\Gamma$. The pair $(Z,\Gamma)$ is dlt and each irreducible component of $\Gamma$ is $\mathbb{Q}$-Cartier. The morphism $h$ is an isomorphism over general points of $S$. The pair $(Z,\Gamma)$ is snc on $h^{-1}(X\setminus S)$.
Then there is a projective bimeromorphic morphism $\varphi\colon Y\rightarrow Z$ with a $\varphi$-ample and $\varphi$-exceptional divisor, such that $Y$ has quotient singularities. Furthermore, $\varphi$ is an isomorphism over $h^{-1}(X\setminus S)$ and over general points of the strict transform of $S$ in $Z$.
@@
${V_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu_{i}}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$${W_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{i}}$${W}$
Since $V$ is normal, the family $(V_{i},G_{i})_{i\in I}$ induces a complex analytic orbispace whose quotient space is $V$. Since $Z\setminus\Gamma$ is isomorphic to an open subset of $X\setminus\Delta$, we deduce from the condition (2) that, if $\mu_{i}\colon V_{i}\rightarrow V$ is the natural morphism, then its divisorial critical locus is contained in $\Gamma$. In addition, the condition (3) implies that $V_{i}$ is smooth on $\mu_{i}^{-1}(V\setminus\Gamma)$.
Let $S_{Z}$ be the strict transform of $S$ in $Z$. Let $\Gamma^{j}$ with j=1,,lj=1,...,l be the irreducible components of $\Gamma$, and let $r^{j}$ be the Cartier index of $\Gamma^{j}$. For every point $z\in h^{-1}(S)$, there is an open neighborhood $T$ of $z$ in $V$ such that each $r^{j}\Gamma^{j}|_{T}$ is defined by some holomorphic function $\sigma^{j}$. Thus we obtain a family $(U_{k})_{k\in K}$ of open subsets of $V$, whose union contains $h^{-1}(S)$, such that $r^{j}\Gamma^{j}|_{U_{k}}$ is connected and defined by some holomorphic function $\sigma^{j}_{k}$. Furthermore, there is an application $\iota\colon K\rightarrow I$ such that $U_{k}$ is contained in $\mu_{\iota(k)}(V_{\iota(k)})$. By convention, we set $\sigma_{k}^{j}=1$ if $\Gamma^{j}|_{U_{k}}$ is empty. Let $U=\bigcup_{k\in K}U_{k}$. Then it is a neighborhood of $h^{-1}(S)$ in $V$.
@@
$U_{k}[\sqrt[Mr^{1}]{\sigma_{k}^{1}},...,\sqrt[Mr^{l}]{\sigma_{k}^{l}}]^{nor}$
and we let $\xi_{k}\colon\overline{U}_{k}\rightarrow U_{k}$ be the natural morphism. Then the $\overline{U}_{k}$’s define a normal complex analytic orbispace with quotient space $U$, by Lemma 2.6 and Lemma 2.11. Let $Q_{k}$ be a connected component of the normalization of $\overline{U}_{k}\times_{Z}V_{\iota(k)}$. It follows that the natural finite morphism $Q_{k}\rightarrow U_{k}$ is Galois of Galois group $H_{k}$, and the collection $(Q_{k},H_{k})_{k\in K}$ induces a normal complex analytic orbispace $\mathfrak{Q}$ with quotient space $U$.
${Q_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\eta_{k}}$${V_{\iota(k)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\overline{U}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\xi_{k}}$${U_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Z}$
We denote by $\eta_{k}\colon Q_{k}\rightarrow\overline{U}_{k}$ the natural morphism. Then by Lemma 2.8, we obtain that $\eta_{k}$ is quasi-étale; furthermore, $Q_{k}$ is smooth in codimension 2 since $V_{\iota(k)}$ is smooth over $U_{k}\setminus\Gamma$.
@@
Let $\widetilde{Q}_{k}\rightarrow Q_{k}$ be the functorial desingularization, which is an isomorphism over the smooth locus of $Q_{k}$. Then there is a natural action of $H_{k}$ on $\widetilde{Q}_{k}$, and we obtain a complex orbifold $\widetilde{\mathfrak{Q}}=(\widetilde{Q}_{k},H_{k})_{k\in K}$ with quotient space $\widetilde{Q}$. In particular, $\widetilde{Q}$ has quotient singularities only. There is an induced projective bimeromorphic morphism $\varphi\colon\widetilde{Q}\rightarrow U$. From the properties of $(Q_{k})_{\operatorname{\textsubscript{\rm sing}}}$ above, we see that the indeterminacy locus of $\varphi^{-1}$ is contained in $h^{-1}(S)$ and is of codimension at least 3 in $Z$. In particular, $\varphi$ is an isomorphism over general points of $S_{Z}$. Furthermore, there is an $\varphi$-exceptional and $\varphi$-ample divisor. Thus by Lemma 2.3, $\varphi$ extends to a projective bimeromorphic morphism $\varphi\colon Y\rightarrow Z$. Since $Z$ is smooth outside $h^{-1}(S)$, we deduce that $Y$ has quotient singularities only. This completes the proof of the lemma. ∎
@@
We will divide the proof into several steps. Let $S_{e}$ with e=1,,re=1,...,r be all the irreducible components of $X_{\operatorname{\textsubscript{\rm sing}}}$ which have codimension 2 in $X$.
@@
Step 1. In this step, we will reduce to the case when $S_{1},...,S_{r}$ are pairwise disjoint. By blowing up at centers contained in at least two of the $S_{e}$’s, we obtain a projective bimeromorphic morphism $h\colon Z\rightarrow X$ such that the strict transforms $h^{-1}_{*}S_{1},...,h^{-1}_{*}S_{r}$ are pairwise disjoint. Let $X^{\prime}\rightarrow Z$ be the projective bimeromorphic morphism constructed in Lemma 2.18, with respect to the collection $\{h^{-1}_{*}S_{1},...,h^{-1}_{*}S_{r}\}$. We observe that the indeterminacy locus of $X\dashrightarrow X^{\prime}$ has codimension at least 3. Up to replacing $X$ by $X^{\prime}$, we may assume that $S_{1},...,S_{r}$ are pairwise disjoint. Furthermore, $X$ has klt singularities.
@@
Step 2. In this step, we will reduce to the situation of Lemma 8.1. Let $X_{e}$ be an open neighborhood of $S_{e}$ in $X$ such that the $X_{e}\cap X_{e^{\prime}}=\emptyset$ if $e\neq e^{\prime}$. We apply Proposition 7.1 to each $X_{e}$. Then we have projective bimeromorphic morphisms $p_{e}\colon X^{\prime}_{e}\rightarrow X_{e}$, which satisfies the properties of Proposition 7.1. Since for each $e$, there is a $p_{e}$-ample and $p_{e}$-exceptional divisor, and since the indeterminacy locus of $(p_{e})^{-1}$ is contained in $S_{e}$ and hence compact, by applying Lemma 2.3 for several times, there is a projective bimeromorphic morphism $p\colon X^{\prime}\rightarrow X$ such that $p=p_{e}$ over $X_{e}$. Let $\Delta$ be the whole reduced $p$-exceptional divisor. Then $p(\Delta)$ has codimension at least 3 in $X$.
@@
In the remainder of the proof, by abuse of notation, we denote $X^{\prime}$ by $X$, and $X_{e}^{\prime}$ by $X_{e}$. If $\Delta_{e}=\Delta|_{X_{e}}$, then $X_{e}$, $\Delta_{e}$ and $S_{e}$ satisfy the conditions (1)-(3) of Lemma 8.1. Furthermore, from now on, it is enough to construct $f\colon Y\rightarrow X$, such that $Y$ has quotient singularities, and $f$ is an isomorphism over general points of each $S_{e}$ and over $X_{\operatorname{\textsubscript{\rm sm}}}\setminus\Delta$. We remark that $X_{\operatorname{\textsubscript{\rm sm}}}\setminus\Delta$ is contained in the snc locus of $(X,\Delta)$.
@@
Step 3. In this step, we will construct a bimeromorphic morphism $h\colon Z\rightarrow X$, which plays the same role for $X_{e}$ as $Z$ for $X$ in Lemma 8.1. We first note that, by the property (3) of Lemma 8.1, $X$ has quotient singularities around $S\setminus\Delta$, where $S=\bigcup_{e=1}^{r}S_{e}$. Thus, we can apply Corollary 3.4 to $(X,\Delta)$, with respect to $S$, and obtain a projective bimeromorphic morphism $\rho\colon X^{\prime}\rightarrow X$ with $X^{\prime}$ normal. By construction, $\rho$ is an isomorphism over the snc locus of $(X,\Delta)$ and over general points of each $S_{e}$; if $\Delta^{\prime}$ is the sum of $\rho^{-1}_{*}\Delta$ and the whole reduced $\rho$-exceptional divisor, then $(X^{\prime},\Delta^{\prime})$ is snc on $\rho^{-1}(X\setminus S)$; furthermore, $\Delta^{\prime}$ is the support of some Cartier divisor, and $X^{\prime}\setminus\Delta^{\prime}$ has quotient singularities. Let $S^{\prime}$ be the strict transform of $S$ in $X^{\prime}$. Let $(Z,\Gamma)\rightarrow(X^{\prime},\Delta^{\prime})$ be a $\mathbb{Q}$-factorial dlt modification as in Lemma 2.19, which is an isomorphism over the snc locus of $(X^{\prime},\Delta^{\prime})$, and over general points of each component of $S^{\prime}$. We denote by $h\colon Z\rightarrow X$ the natural morphism. Then we have the following properties. The divisor $\Gamma$ is the sum of $h^{-1}_{*}\Delta$ and the whole reduced $h$-exceptional divisor. The $h$-exceptional locus is contained in $\Gamma$. Each irreducible component of $\Gamma$ is $\mathbb{Q}$-Cartier. The morphism $h$ is an isomorphism over the snc locus of $(X,\Delta)$ and over general points of each component of $S$. The pair $(Z,\Gamma)$ is snc on $h^{-1}(X\setminus S)$.